If two consecutive even numbers are such that 5/6 of the reciprocal of one is
equal to the reciprocal of the other. Find the two even numbers.

Answer by mananth(16946)

If two consecutive even numbers are such that 5/6 of the reciprocal of one is
equal to the reciprocal of the other. Find the two even numbers.
...
let the numbers be x, x+2
..
5/6x= 1/(x+2)
cross multiply
5(x+2)=6x
5x+10=6x
6x-5x=10
x=10
...
10,12 are the numbers